# Progenitor diversity in the accreted stellar halos of Milky Way-like galaxies

### v1

@@ 4 Conclusions
We have explored the diversity of the present-day accreted stellar halos of Milky Way mass galaxies by quantifying the fraction of mass contributed by different progenitors as a function of galactocentric radius, in three different sets of cosmological simulations: BJ05, Aquarius and Auriga. These models cover a range around the likely virial mass of the Milky Way (see Fig. 1) and use very different methods for populating progenitor dark matter halos with stars. Each model has been used extensively in the literature to predict and interpret observations of the real Milky Way. We find systematic differences in the composition of accreted stellar halos predicted by these models, notwithstanding the galaxy-to-galaxy differences within any given model owing to the stochastic nature of progenitor masses, accretion times and orbits. Specifically, we find:
@@
    The BJ05 halos have a much larger number of significant progenitors than the other two simulations, both globally (Figs. 4 and  5) and as a function of radius (Fig. 7). This higher diversity, reflecting assembly from a relatively larger number of lower-mass progenitors, is most likely due to assumptions made in the generation of the BJ05 models. For example, at a fixed progenitor virial mass, the BJ05 models predict stellar masses for the most massive progenitors up to an order of magnitude lower than in Auriga (Fig. 2).
@@
    The Auriga simulations are broadly similar to the Aquarius models in the region of the accreted stellar halo probed by current observations, but they predict increasing diversity at a larger radius, whereas the Aquarius simulations predict a peak in diversity around $\sim 50$ kpc. These differences may again be a consequence of systematic differences in the star formation histories of progenitors (the Aquarius progenitors form stars less efficiently than those in Auriga, producing less massive accreted stellar halos overall), combined with differences in the tidal disruption process between the hydrodynamical and particle-tagging method (for a detailed discussion see Cooper et al., 2017).
@@
    Essentially none of the simulations reproduce the combination of very low diversity and early assembly implied by the Milky Way data;
@@
    The majority of systems with diversity as low as that implied for the Milky Way, in the currently-observed region of the halo, are examples of relatively recent accretion of a massive progenitor. This scenario is effectively ruled out for the Milky Way on chemical and dynamical grounds;
@@
    There are several simulated galaxies that have broadly similar progenitor mass deposition profiles (‘strata’) and corresponding progenitor infall times to those inferred for Milky Way, in particular in Auriga. However, in these cases the number of dominant progenitors is higher than the $\sim 2$ currently suggested for the Milky Way. On this basis, the simulations predict that other significant progenitor debris systems, closely resembling the chemo-dynamical structure of GS/E, await identification in the Milky Way.
@@
An important caveat for the work presented here is that the Naidu et al. (2020) result refers to a specific, nontrivial survey geometry, whereas (for simplicity and generality) our analysis of the simulations refers to a complete spherical volume. The above conclusions could be changed if the Naidu et al. (2020) result is not representative of the entirety of the Milky Way’s accreted stellar halo in the same radial range, for example if accreted debris on circular orbits in the plane of the Milky Way disk makes a significant contribution (e.g. Gómez et al., 2017).
@@
As we describe in the introduction, most work on the diversity of accreted stellar halos predicts that 1similar-toabsent1$\sim$ 1∼ 110101010 progenitors dominate the total accreted mass. This corresponds well to currently available evidence for the diversity of the real Milky Way halo (and the halos of other Milky Way-like galaxies, e.g. Monachesi et al., 2016a; Gozman et al., 2023). However, the composition of the Milky Way’s stellar halo may eventually be constrained in much greater detail than that of any other galaxy. In that context, the difference between one and ten significant progenitors is still substantial. It is therefore worthwhile to consider the implications of the apparently low diversity of the Milky Way’s stellar halo suggested by current data – in particular, whether that finding is indicative of a particular subset of possible assembly histories.
@@
A broad range of simulations of simulated systems with comparable virial mass, using different techniques, can provide useful context and intuition when interpreting reconstructions of the Milky Way’s history. In that regard, we believe it is notable that the simulations we have examined here show many cases of galaxies in the virial mass range of the Milky Way with a larger number of significant stellar halo progenitors, by $N_{\mathrm{sig}}$ or other measures (e.g. Monachesi et al., 2019). Moreover, halos with high diversity overall may appear to be dominated by only one or two progenitors if observed over a restricted range of radius (and vice versa). As noted by Deason et al. (2016), pre-selecting galaxy assembly histories, for example to exclude major mergers, biases any conclusions regarding what kind of stellar halo assembly scenarios (and hence diversities) are typical for halos of a particular virial mass.
@@
Deason et al. (2016) conclude that the Milky Way is atypical in hosting several particularly massive surviving satellites (in particular the LMC and SMC; see also e.g. Cautun et al., 2014) while also having an apparently ‘quiescent’ accretion history, on the evidence of its stable disk and low-mass, metal-poor stellar halo. They note that this situation may be transitory; the disruption of the LMC and the SMC will, in this and other respects ‘restore the Milky Way to the mean’ (Cautun et al., 2019). A stellar halo dominated by two recently-accreted progenitors, with somewhat correlated orbits would resemble the example of Au08 presented above, and hence would be more typical of the low-diversity galaxies we find in the simulations.
@@
Our results suggest that predictions of diversity from simulations are strongly model-dependent. Clearly it would be valuable, and relatively straightforward, to make similar predictions with the many other Milky Way-analogue zoom simulations now available. Our work suggests that such predictions can be usefully compared to ongoing and future Milky Way surveys, including DESI MWS (Cooper et al., 2023), 4MOST (de Jong et al., 2019), WEAVE (Jin et al., 2023), SDSS-V (Kollmeier et al., 2017) and PFS (Greene et al., 2022). These surveys will further improve the constraints on halo pioneered by Naidu et al. (2020), in particular by extending the radial range studied and the number of stars observed.
@@
PSY and APC are supported by a Yushan Fellowship awarded by the Taiwanese Ministry of Education, MOE-113-YSFMS-0002-001-P2. APC acknowledges support from Taiwan’s National Science and Technology Council under grants 109-2112-M-007-011-MY3, 112-2112-M-007-017 and 113-2112-M-007-009. AM gratefully acknowledges support by FONDECYT Regular grant 1212046 and by the ANID BASAL project FB210003, as well as funding from the Max Planck Society through a “PartnerGroup” grant and the HORIZON-MSCA-2021-SE-01 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement number 101086388. RG is supported by an STFC Ernest Rutherford Fellowship (ST/W003643/1). This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University. This equipment was funded by the Ministry of Education of Taiwan, the National Science and Technology Council of Taiwan, and National Tsing Hua University. We have used simulations from the Auriga Project public data release (Grand et al., 2024) available at https://wwwmpa.mpa-garching.mpg.de/auriga/data.
@@
We provide the Cooper et al. (2010) star particle data for Aquarius at https://doi.org/10.5281/zenodo.13888986: this is the first public release of those data. A description of the data and related scripts are available at https://github.com/nthu-ga/aquarius-halos, as described in the appendix. Our progenitor IDs for the stellar particles in the Auriga data release are provided at https://doi.org/10.5281/zenodo.13943963.
@@ Appendix A Stellar Halo Data
We make the Aquarius simulation data of Cooper et al. (2010) publicly available in HDF5 format at https://doi.org/10.5281/zenodo.13888986. The data model for these files is described in Table 2 and further documented at https://github.com/nthu-ga/aquarius-halos. The dataset names and the file header follow the conventions of the Gadget file format and should be compatible with most routines designed to read that format. At the above URL, we also provide scripts to convert the Bullock & Johnston (2005) simulation outputs to that format, and Progenitor ID labels for accreted star particles in the L3 Auriga halos. We plan to update these data products with extra properties (such as progenitor infall time, metallicity and age) and to provide other related data products in association with future publications.
The rows of each dataset in these files correspond to ‘star particles’ in the simulations. Only accreted particles are included. The Cartesian coordinate frame is that of the centre of the galactic potential. The field ProgenitorID identifies the last branch of the subhalo merger tree to which a star particle was bound before it became bound to the dark matter halo of the central galaxy. These labels should be treated as arbitrary integers. Particles currently bound to satellite subhalos can be identified by SubhaloNr $>0$. For further details, see the documentation URL above.

### v2

@@ 4 Conclusions
We have explored the diversity of the present-day accreted stellar halos of Milky Way mass galaxies by quantifying the fraction of mass contributed by different progenitors as a function of galactocentric radius, in three different sets of cosmological simulations: BJ05, Aquarius, and Auriga. These models cover a range around the likely virial mass of the Milky Way (see Figure 1) and use very different methods for populating progenitor dark matter halos with stars. Each model has been used extensively in the literature to predict and interpret observations of the real Milky Way. We find systematic differences in the composition of accreted stellar halos predicted by these models, notwithstanding the galaxy-to-galaxy differences within any given model owing to the stochastic nature of progenitor masses, accretion times, and orbits. Specifically, we find:
@@
    The BJ05 halos have a much larger number of significant progenitors than the other two simulations, both globally (Figs. 4 and  5) and as a function of radius (Figure 7). This higher diversity, reflecting assembly from a relatively larger number of lower-mass progenitors, is most likely due to assumptions made in the generation of the BJ05 models. For example, at a fixed progenitor virial mass, the BJ05 models predict stellar masses for the most massive progenitors up to an order of magnitude lower than in Auriga (Figure 2).
@@
    The Auriga simulations are broadly similar to the Aquarius models in the region of the accreted stellar halo probed by current observations, but they predict increasing diversity at a larger radius, whereas the Aquarius simulations predict a peak in diversity around $\sim 50$ kpc. These differences may again be a consequence of systematic differences in the star formation histories of progenitors (the Aquarius progenitors form stars less efficiently than those in Auriga, producing less massive accreted stellar halos overall), combined with differences in the tidal disruption process between the hydrodynamical and particle-tagging method (for a detailed discussion, see Cooper et al., 2017).
@@
    Essentially none of the simulations reproduce the combination of very low diversity and early assembly implied by the Milky Way data.
@@
    The majority of systems with diversity as low as that implied for the Milky Way, in the currently observed region of the halo, are examples of relatively recent accretion of a massive progenitor. This motivates further consideration of the accretion times of GS/E and Sgr.
@@
    There are several simulated galaxies that have broadly similar progenitor mass deposition profiles (‘strata’) and corresponding progenitor infall times to those inferred for Milky Way, in particular in Auriga. However, in these cases, the number of dominant progenitors is higher than the $\sim 2$ currently suggested for the Milky Way. On this basis, the simulations predict that other significant progenitor debris systems, closely resembling the chemodynamical structure of GS/E, await identification in the Milky Way.
@@
An important caveat for the work presented here is that the Naidu et al. (2020) result refers to a specific, nontrivial survey geometry, whereas (for simplicity and generality) our analysis of the simulations refers to a complete spherical volume. The above conclusions could be changed if the Naidu et al. (2020) result is not representative of the entirety of the Milky Way’s accreted stellar halo in the same radial range—for example, if accreted debris on circular orbits in the plane of the Milky Way disk makes a significant contribution (e.g. Gómez et al., 2017). The apparent discrepancy with the simulations would also be reduced if the total stellar mass of GS/E or Sgr was found to be an overestimate (e.g. Lane et al., 2023).
@@
As we describe in the introduction, most work on the diversity of accreted stellar halos predicts that 1similar-toabsent1$\sim$ 1∼ 110101010 progenitors dominate the total accreted mass. This corresponds well to currently available evidence for the diversity of the real Milky Way halo (and the halos of other Milky Way–like galaxies, e.g., Monachesi et al., 2016a; Gozman et al., 2023). However, the composition of the Milky Way’s stellar halo may eventually be constrained in much greater detail than that of any other galaxy. In that context, the difference between 1 and 10 significant progenitors is still substantial. It is therefore worthwhile to consider the implications of the apparently low diversity of the Milky Way’s stellar halo suggested by current data – in particular, whether that finding is indicative of a particular subset of possible assembly histories.
@@
A broad range of simulations of simulated systems with comparable virial mass, using different techniques, can provide useful context and intuition when interpreting reconstructions of the Milky Way’s history. In that regard, we believe it is notable that the simulations we have examined here show many cases of galaxies in the virial mass range of the Milky Way with a larger number of significant stellar halo progenitors, by $N_{\mathrm{sig}}$ or other measures (e.g. Monachesi et al., 2019). Moreover, halos with high diversity overall may appear to be dominated by only one or two progenitors if observed over a restricted range of radius (and vice versa). As noted by Deason et al. (2016), preselecting galaxy assembly histories, for example to exclude major mergers, biases any conclusions regarding what kind of stellar halo assembly scenarios (and hence diversities) are typical for halos of a particular virial mass.
@@
Deason et al. (2016) conclude that the Milky Way is atypical in hosting several particularly massive surviving satellites (in particular, the LMC and SMC; see also, e.g., Cautun et al., 2014) while also having an apparently ‘quiescent’ accretion history, on the evidence of its stable disk and low-mass, metal-poor stellar halo. They note that this situation may be transitory; the disruption of the LMC and the SMC will, in this and other respects ‘restore the Milky Way to the mean’ (Cautun et al., 2019). A stellar halo dominated by two recently accreted progenitors with somewhat correlated orbits would resemble the example of Au08 presented above, and hence would be more typical of the low-diversity galaxies we find in the simulations.
@@
Our results suggest that predictions of diversity from simulations are strongly model-dependent. Clearly, it would be valuable, and relatively straightforward, to make similar predictions with the many other Milky Way–analog zoom simulations now available. Our work suggests that such predictions can be usefully compared to ongoing and future Milky Way surveys, including DESI MWS (Cooper et al., 2023), 4MOST (de Jong et al., 2019), WEAVE (Jin et al., 2023), SDSS-V (Kollmeier et al., 2017), and PFS (Greene et al., 2022). These surveys will further improve the constraints on halo pioneered by Naidu et al. (2020), in particular by extending the radial range studied and the number of stars observed.
@@
We thank the anonymous reviewer for a thorough and constructive report, which greatly improved the clarity of the paper and brought important related work to our attention. PSY and APC are supported by a Yushan Fellowship awarded by the Taiwanese Ministry of Education (MOE), MOE-113-YSFMS-0002-001-P2. APC acknowledges support from Taiwan’s National Science and Technology Council (NSTC) under grants 109-2112-M-007-011-MY3, 112-2112-M-007-017 and 113-2112-M-007-009. AM gratefully acknowledges support by FONDECYT Regular grant 1212046 and by the ANID BASAL project FB210003, as well as funding from the Max Planck Society through a ‘PartnerGroup’ grant and the HORIZON-MSCA-2021-SE-01 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement number 101086388. RG is supported by an STFC Ernest Rutherford Fellowship (ST/W003643/1). This work used high-performance computing facilities operated by the Center for Informatics and Computation in Astronomy (CICA) at National Tsing Hua University (NTHU). This equipment was funded by MOE, NSTC, and NTHU. We have used the Auriga Project data release (Grand et al., 2024) available at https://wwwmpa.mpa-garching.mpg.de/auriga/data.
@@
We provide the Cooper et al. (2010) star particle data for Aquarius at https://doi.org/10.5281/zenodo.13888986: this is the first public release of those data. A description of these data and related scripts are available at https://github.com/nthu-ga/aquarius-halos, as described in the appendix. Our progenitor IDs for the stellar particles in the Auriga data release are provided at https://doi.org/10.5281/zenodo.13943963.
@@ Appendix A Stellar Halo Data
We make the Aquarius simulation data of Cooper et al. (2010) publicly available in HDF5 format at https://doi.org/10.5281/zenodo.13888986. The data model for these files is described in Table 2 and further documented at https://github.com/nthu-ga/aquarius-halos. The dataset names and the file header follow the conventions of the Gadget file format and should be compatible with most routines designed to read that format. At the above URL, we also provide scripts to convert the Bullock & Johnston (2005) simulation outputs to the same format we use for Aquarius, along with progenitor ID labels for accreted star particles in the L3 Auriga halos. We plan to update these data products with extra properties (such as progenitor infall time, metallicity, and age) and to provide other related data products in association with future publications.
The rows of each dataset in these files correspond to ‘star particles’ in the simulations. Only accreted particles are included. The Cartesian coordinate frame is that of the center of the galactic potential. The field ProgenitorID identifies the last branch of the subhalo merger tree to which a star particle was bound before it became bound to the dark matter halo of the central galaxy. These labels should be treated as arbitrary integers. Particles currently bound to satellite subhalos can be identified by SubhaloNr $>0$. For further details, see the documentation URL above.